OpenStudy (imsleeziboii):

2 questions for medal!!!!!!!!!!!!

rebeccaxhawaii (rebeccaxhawaii):

one per post but what is your question

OpenStudy (imsleeziboii):

https://media.education2020.com/evresources/2003-08-05-00-00_files/i0020000.jpg Write an equation for the line by using ordered pairs from the line to find a function rule. A. y = 3x + 3 B. y = 2x – 2 C. y = 2x + 3 D. y = 3x

rebeccaxhawaii (rebeccaxhawaii):

what is the slope of the line

rebeccaxhawaii (rebeccaxhawaii):

pick two points

OpenStudy (imsleeziboii):

I dont know how to find slope

OpenStudy (imsleeziboii):

-2,-5

rebeccaxhawaii (rebeccaxhawaii):

pick two points

OpenStudy (imsleeziboii):

-2,-5

rebeccaxhawaii (rebeccaxhawaii):

those arent points

OpenStudy (imsleeziboii):

what do you mean by points then

rebeccaxhawaii (rebeccaxhawaii):

there the line meets the graph

OpenStudy (imsleeziboii):

3,-2?

rebeccaxhawaii (rebeccaxhawaii):



rebeccaxhawaii (rebeccaxhawaii):

that circle is where a point is because it meets the line and graph

rebeccaxhawaii (rebeccaxhawaii):

so pick two points on your graph

OpenStudy (imsleeziboii):

2

OpenStudy (imsleeziboii):

-2

rebeccaxhawaii (rebeccaxhawaii):

no not where it hits the number

rebeccaxhawaii (rebeccaxhawaii):

it seems you don't know a clue about this you need to review your lesson

OpenStudy (imsleeziboii):

i cant review my lesson. lol f.o.

rebeccaxhawaii (rebeccaxhawaii):



rebeccaxhawaii (rebeccaxhawaii):

pick a point that the line lands on one of these

OpenStudy (imsleeziboii):

i dont get it